A jar contains 5-gram bolts and 10-gram bolts. The contents of the jar weigh 2.35kg. If there are 300 bolts altogether, how many are there of each kind?
Answer by checkley71(8403)

5x+10(300-x)=2350
5x+3000-10x=2350
-5x=2350-3000
-5x=-650
x=-650/-5
x=130 THE NUMBER OF5 gram bolts in the jar.
300-130=170 THE NUMBER OF 10 gram bolts in the jar.
proof
5*130+10*170=2350
650+1700=2350
2350=2350